OpenStudy (anonymous):

Solve for x: .95=[(e^x)-1]/x Step by step would be appreciated

OpenStudy (jdoe0001):

\(\bf 0.95=\cfrac{e^x-1}{x}\qquad or \qquad 0.95=\cfrac{e^{x-1}}{x}\quad ?\)

OpenStudy (zzr0ck3r):

lets hope the first one, else im not sure anyone here can do it.

OpenStudy (jdoe0001):

heh

OpenStudy (anonymous):

yeah sorry it's the first one

OpenStudy (shamil98):

\[0.95 = \frac {e^x - 1} {x}\] \[\ 0.95x = e^x -1\] \[\ 0.95x + 1 = e^x \]

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

"Solve for x:" =)

OpenStudy (shamil98):

well yeah, you would just take the log of both sides from there

OpenStudy (jdoe0001):

well.. ahemm wouldn't that land us at ln(0.95x+1) = x ?

OpenStudy (shamil98):

\[\ \ln(.95x + 1) = \ln e^x\] \[\ \ln (.95x + 1) = x \] hehe

OpenStudy (anonymous):

that's where i got stuck

OpenStudy (jdoe0001):

hehe

OpenStudy (shamil98):

hmm

OpenStudy (shamil98):

wolfram gave a weird answer